# PyAtoms: An interactive tool for rapidly simulating atomic scanning tunneling microscopy images of 2D materials, moiré systems and superlattices

### v2

@@ 2.4 Homogeneous strain
Strain is a powerful experimental method to tune electronic and optical properties of materials. When a crystal lattice is strained, the atomic bonds in real-space become distorted, consequently shifting the reciprocal lattice vectors. This can drastically alter a material’s electronic properties by distorting the shape of the Fermi surface and the Brillouin zone boundaries, affecting the vibrational properties, and even opening of band gaps in the band structure [42].
To leading order in the strain tensor, $\boldsymbol{\epsilon}$, the presence of strain affects the direct lattice vectors via
$\mathbf{v}=(\mathbb{I}+\boldsymbol{\epsilon})\mathbf{v}_{0},$ (16)
where $\mathbf{v}_{0}$ is the un-strained direct lattice vector, $\mathbf{v}$ is the strained direct lattice vector, $\mathbb{I}$ is the $2\times 2$ identity matrix, and $\boldsymbol{\epsilon}$ is the two-dimensional strain tensor given by
$\boldsymbol{\epsilon}=\begin{pmatrix}\epsilon_{xx}&\epsilon_{xy}\\ \epsilon_{yx}&\epsilon_{yy}\end{pmatrix}.$ (17)
In PyAtoms, the strain $\hat{x}$-axis is directed along the first direct lattice vector $\mathbf{a}_{1}=a\hat{x}$. The resulting strain-distorted reciprocal lattice vectors are given by (to leading order in $\boldsymbol{\epsilon}$),
@@ 2.5 Low-pass filtering
3 PyAtoms Overview
@@
(a,b,c) Lattice 1,2,3: In the Lattices section, users select the properties for each constituent layer. In the Parameters tab users select: the lattice symmetry (triangular/hexagonal or square), which sets the primitive lattice vectors; the lattice constant, in nanometers; and twist angle ($\theta_{12},\theta_{23}$) for two- or three-layer systems. For triangular/hexagonal systems, the Sublattices tab allows users to select: the lattice site at the image origin (hollow, A, or B) and the amplitude of each sublattice ($\alpha_{i},\beta_{i}$) (see Fig. 1). In the Strain tab, users have the option to apply a 2D strain tensor (Eq. 17), where the $x$-axis of strain tensor for $i$th layer is along the direction of the $i$th layer’s first reciprocal lattice vector, $\mathbb{g}_{1}^{i}$ (Eqs. 3, 4).
@@
(f) Low pass filtering: For multilayer simulations, users have the option input a value for sigma, the radius of Gaussian mask in real space (pixels) units. The full-width at half-max (FWHM) of the gaussian filter is shown as a white circle in the bottom left corner of the real space image. Likewise, the FWHM of the gaussian filter in reciprocal space is shown as a red circle in the FFT (Fig. 4(d,f)). The radius of the gaussian mask in nanometers is shown in real-time in the text box.
@@
where $N_{pix}$ is the number of pixels in one line, $L$ is the length of the image in nanometers, $v_{t}$ is the velocity (in nm/s) of the tip in one direction and the factor of 2 takes into account both the left and right fast-scan directions. This section is useful in estimating the time to record a “closed-feedback” $dI/dV_{b}$ map where the LDOS($eV_{b}$) is recorded at a single energy as the tip slowly scans the surface.
@@
The user inputs the time per spectra, which is the total time in seconds (including overhead) to record a single spectrum, $T_{spec}$, i.e., one $dI/dV_{b}$ sweep. This section is useful in estimating the time to record an “open-feedback” LDOS($eV_{b}$) map where a $dI/dV_{b}(V_{b})$ spectra is recorded at every pixel position. Together with the imaging window (Fig. 3(g)) provided by PyAtoms, this estimator is invaluable in ensuring that time-consuming QPI measurements will have the requisite reciprocal space resolution and spatial extent while avoiding aliasing effects.
@@ 4.1 Research
With its simple and real-time GUI interface, PyAtoms is a powerful research tool during STM data collection as well as for data analysis. As shown in Fig. 2, PyAtoms can be used side-by-side during STM measurements in order to rapidly simulate experimental images and determine properties such as the constituent lattice constants and moiré twist angles between layers. Below we describe two examples of PyAtoms research capabilities.
@@
In practice, the strain tuning capabilities of PyAtoms can be used to simulate and determine the amount of true, physical strain in STM images or the apparent strain due to scan artifacts such as tip drift. Figure 4 demonstrates the dramatic effect that strain can have on the moiré lattice and how PyAtoms’ can be used to determine this strain. Fig. 4(a) is a simulation of twisted trilayer graphene ($\theta_{12}=1.42^{\circ},\theta_{23}=-1.88^{\circ}$) moiré quasicrystal [43] that features two moiré lattices. The atomic Bragg peaks for each graphene layer can be seen in the Fourier transform (Fig. 4(b)). The complex spatial structure of the moiré -of- moiré pattern (Fig. 4(c)) can be better seen after low-pass filtering to keep only the long wavelength periodicities (Fig. 4(d)). The presence of even miniscule strain in the top-most graphene layer ($\epsilon_{xx}=0.25$%) significantly alters the larger moiré pattern in both real (Fig. 4(e)) and reciprocal space (4(f)). For researchers studying moiré multi-layers, PyAtoms can thus be used side-by-side during measurements to quickly determine the local strain in an area so that users can make informed decisions whether to seek other, less-strained areas.
[…]
4.2 PyAtoms in education
PyAtoms’ speed and user-friendly interface makes it a unique interactive educational tool for teaching fundamental concepts in Fourier analysis and crystal structure. One of us (C.G.) has used PyAtoms as part of lecture demonstrations in two different undergraduate courses: sophomore-level “Mathematical Methods for Physicists” and senior-level “Introduction to Solid State Physics.” In the maths course, PyAtoms’ real-time imaging of the real- and reciprocal-space structures was used to convey concepts in 2D Fourier analysis, including the real- and reciprocal representation of plane waves and the concept of aliasing. Hiding half of the screen, students would be presented with either a real-space image and asked to determine the Fourier-space structure, or vice versa. In the solid state course, PyAtoms has been used in lecture to reinforce concepts in crystal structure, such as determining the primitive direct and reciprocal lattice vectors. It has also been used to connect students to modern concepts in 2D materials, such as moiré systems.
To illustrate the value of PyAtoms to solid state physics education, consider the problem of determining both the lattice and commensurate superlattice parameters $(n,\phi)$ for the system shown in Fig. 6. Here, students are provided the raw text data generated by PyAtoms and are tasked with importing and visualizing the data and calculating the 2D Fourier transform.
Figure 6: Square superlattice. (a) PyAtoms simulation of a square superlattice $(\sqrt{5}\times\sqrt{5})R26.57^{\circ}$. Inset: Zoom view of the smaller square lattice ($a_{1}=3$, yellow square) and the ($\sqrt{5}\times\sqrt{5})R26.57^{\circ}$ superlattice ($a_{2}=3\times\sqrt{5}\approx 6.71$, blue square). (b) FFT of the image in (a). Yellow circles denote the atomic Bragg peaks; light blue circles denote the 1st order superlattice peaks and light purple the satellite superlattice peaks.
Question: The data contains both a lattice and a superlattice. What are the primitive lattice vectors for each lattice? What values of $(n,\phi)$ describe the superlattice?
Solution: Real and FFT analysis of Fig. 6 shows that the smaller, atomic lattice is a square lattice with primitive lattice vectors given by $\mathbf{a_{1}}=a\hat{x}$ and $\mathbf{a_{2}}=a\hat{y}$, where $a=3$. By visual inspection of Fig. 6 (inset), one sees that the square superlattice vectors are given by $\mathbf{b_{1}}=2\mathbf{a_{1}}+\mathbf{a_{2}}$ and $\mathbf{b}_{2}=-\mathbf{a}_{1}+2\mathbf{a}_{2}$. The objective is to first determine $n$, which is the magnitude of $\mathbf{b_{1}}$:
$\mathbf{b_{1}}$ $=2a\hat{x}+a\hat{y}$
$\Rightarrow|\mathbf{b_{1}}|$ $=\sqrt{2^{2}+1^{2}}a=\sqrt{5}a.$
The orientation of the superlattice with respect to the original, $\phi$, is then given by simple geometry:
$\tan{\phi}$ $=\frac{b_{1x}}{b_{1y}}=\frac{1}{2}\Rightarrow\phi=\arctan{\frac{1}{2}}\approx 26.57^{\circ}$
$\Rightarrow n$ $=\sqrt{5},\phi=26.57^{\circ}$
Problem sets containing this $(\sqrt{5}\times\sqrt{5})R26.57^{\circ}$ square system and other superlattices (see Fig. 5) have been assigned in senior-level solid state physics courses at UCLA by C.G. for several years with great success, as determined by post-course evaluations. Here, PyAtoms was used to generate the raw data which students are directed to import, visualize (real and Fourier transformed images) as if it were actual data recorded in an STM laboratory. This example showcases the educational value of PyAtoms, highlighting its practical application in reinforcing theoretical concepts concerning primitive lattice vectors and determining crystal structure as well as inspiring students to develop key programming and data analysis skills. The ease with which complex systems can be simulated, such as multi-layer moiré systems (Fig. 4), offers educators the opportunity to connect their students to concepts in current condensed matter physics research.
5 Outlook
We have presented PyAtoms, an interactive open-source Python-based software for the real-time simulation of STM images of 2D materials. With its quick, intuitive, and robust simulation capabilities, we envision PyAtoms as valuable companion during both STM measurements and STM data analysis. PyAtoms can also be used as an educational tool to train new STM researchers as well as in the classroom when covering concepts in Fourier analysis and solid state physics.
@@ 5 Outlook
Possible future developments include offering other methods for approximating images of multilayer systems and allowing generic 2D lattices to be generated beyond square and triangular. Similarly, for the CDW/BDW simulations, control of the phase in real space would allow easier comparisons to real STM data [6, 31, 32]. A current weakness in PyAtoms, is that comparison to experiment is largely qualitative. However, apart from its GUI, the open-source code that underlies PyAtoms can be called separately (through a Python terminal, for instance) and can be used, for example, to simulate moiré and superlattice systems beyond three layers. In principle, this ability to call on PyAtoms’ base functions offers the intriguing possibility to determine the simulation parameters (lattice constant, A- and B-sublattice amplitude, strain) through a least squares fitting to experimental STM data. This latter capability would elevate PyAtoms to a useful quantitative tool in the study of 2D materials.
Acknowledgments
@@
The authors thank M. Mandigo-Stoba, K.-Y. Wey, and S. Nekarae for providing STM images, F. Joucken for carefully explaining the “Log” method of STM simulation in reference [38], and J. Berger, C. Irving for technical assistance with STM assembly. A.G.P acknowledges financial support from the National Science Foundation (NSF), the UCLA office of the Dean of Physical Sciences through the UCLA Physics Bridge Program and the UCLA Graduate Division through a Eugene Cota-Robles Fellowship. C.G. acknowledges teaching relief support from the UCLA offices of the Executive Vice Chancellor and Provost, Vice Chancellor for Research and Creative Activities, and the Vice Chancellor for Academic Personnel and partial support from the UCLA Hellman Society of Fellows.
@@
C.G. created the original PyAtoms base codes (hexatoms, squareatoms, moirelattice) in Matlab. A.G.P. translated the base codes from Matlab to Python and created the GUI interface with input from C.G. Both A.G.P. and C.G. contributed to PyAtoms’ later updates. C.G. wrote the manuscript and created the figures with input from A.G.P.

### v3

@@ II.3 Superlattices and multi-layered systems
We note other recent analytical models for simulating moiré and/or multilayered STM images. Le Ster et al le2024moire have presented the moiré plane wave expansion (MPWE) model which produces a total STM image that includes interactions between layers by weighing the contributions of three terms: (i) a non-interacting substrate layer, $z(\mathbf{r})$ composed of plane waves with wavevectors $\mathbf{G}_{mn}$; (ii) a non-interacting second layer, $z^{\prime}(\mathbf{r})$ composed of plane waves with wavevectors $\mathbf{G^{\prime}}_{pq}$; and a moiré term, $z_{M}(\mathbf{r})$, which takes into account interactions by being composed of plane waves with wavevectors, $\mathbf{K}_{mnpq}=\mathbf{G^{\prime}}_{pq}-\mathbf{G}_{mn}$, which are all possible linear combinations of wavevectors of the constituent layers le2019moire. Pásztor et al pasztor2024delusive have presented a similar model which, for the purpose of investigating artifacts in geometric phase analyses, consists of summing (without weights) plane waves with all first-order linear combinations of wavevectors of the constituent lattices. Future versions of PyAtoms can easily allow these and even more models. We stress that, since the goal of PyAtoms is to rapidly simulate ideal moiré and other super-lattices, the simple sums of plane waves cannot account for lattice relaxation that can occur in real systems woods_commensurateincommensurate_2014.
II.4 Homogeneous strain
Lattice strain can strongly affect the electronic and optical properties of materials. When a crystal lattice is strained, the atomic bonds in real-space become distorted, consequently shifting the reciprocal lattice vectors. This can distort the shape of the Fermi surface and the Brillouin zone boundaries, affecting the vibrational properties, and even the opening of band gaps in the band structure pereira_tight-binding_2009.
To leading order in the strain tensor, $\bm{\epsilon}$, the presence of strain affects the direct lattice vectors via
$\mathbf{v}=(\mathbb{I}+\bm{\epsilon})\mathbf{v}_{0},$ (16)
where $\mathbf{v}_{0}$ is the un-strained direct lattice vector, $\mathbf{v}$ is the strained direct lattice vector, $\mathbb{I}$ is the $2\times 2$ identity matrix, and $\bm{\epsilon}$ is the two-dimensional strain tensor given by
$\bm{\epsilon}=\begin{pmatrix}\epsilon_{xx}&\epsilon_{xy}\\ \epsilon_{yx}&\epsilon_{yy}\end{pmatrix}$. (17)
In PyAtoms, the strain $\hat{x}-$axis is directed along the first direct lattice vector $\mathbf{a}_{1}=a\hat{x}$. The resulting strain-distorted reciprocal lattice vectors are given by (to leading order in $\bm{\epsilon}$),
@@ II.5 Low-pass filtering
Figure 5: Setting up and optimizing QPI images with PyAtoms. (a) Simulated FT-STM ($624\times 624$ pixels, 60 nm) of NbSe2 with three-fold $(4\times 1)R0^{\circ}$ CDW domains. (b) Experimental STM topography image of NbSe2 at $T=4.7$ K with strain-induced $(4\times 1)R0^{\circ}$ disordered CDW domains ($V_{b}=1.5$ mV, $I_{t}=3$ pA). Image has optimized scan parameters from (a). Image acquisition time $\approx 90$ min. (c) FT-STM of the topography in (b) showing additional features corresponding to QPI patterns seen previously gao2018atomic.
III PyAtoms Overview
@@
(a, b, c) Lattice 1, 2, 3: In the Lattices section, users select the properties for each constituent layer. In the Parameters tab users select: the lattice symmetry (triangular/hexagonal or square), which sets the primitive lattice vectors; the lattice constant, in nanometers; and the twist angle ($\theta_{12},\theta_{23}$) for two- or three-layer systems. For triangular/hexagonal systems, the Sublattices tab allows users to select: the lattice site at the image origin (hollow, A, or B) and the amplitude of each sublattice ($\alpha_{i},\beta_{i}$) (see Fig. 1). In the Strain tab, users have the option to apply a 2D strain tensor (Eq. 17), where the $x$-axis of strain tensor for $i$th layer is along the direction of the $i$th layer’s first reciprocal lattice vector, $\mathbb{g}_{1}^{i}$ (Eqs. 3, 4).
@@
Figure 6: Simulating bias-dependent $\bm{(2\times 2)R0^{\circ}}$ CDWs in a transition metal dichalgogenide. (a) PyAtoms simulation of NbSe2 atomic lattice using sublattice parameters $\alpha_{at}=0.7,\beta_{at}=0.3$ and (b) the $(2\times 2)R0^{\circ}$ CDW lattice using sublattice parameters $\alpha_{cdw}=0.7,\beta_{cdw}=0.3$. (c) Simulation of the combined atomic and CDW lattices using the ‘Simple’ option ($\eta=0.2$). (d)-(f) Experimental STM topographs of uniaxial strain-induced commensurate $(2\times 2)R0^{\circ}$ CDW (diamond) in NbSe2 recorded at $T=4.7$ K and $I_{t}$ = 200 pA. Voltage biases, $V_{b}$, are (d) 1.3-1.3 V, (e)1.4-1.4 V, and (f) 1.6-1.6 V. Corresponding PyAtoms simulations are shown in (g)-(i) using the ‘Simple’ option and CDW sublattice parameters (g) $\alpha_{cdw}=0.77,\beta_{cdw}=0.23$, (h) $\alpha_{cdw}=0.55,\beta_{cdw}=0.45$, (i) $\alpha_{cdw}=0.98,\beta_{cdw}=0.02$. All scale bars are 1 nm.
(f) Low pass filtering: For multilayer simulations, users have the option to input a value for sigma, the radius of Gaussian mask in real space (pixels) units. The full-width at half-max (FWHM) of the Gaussian filter is shown as a white circle in the bottom left corner of the real space image. Likewise, the FWHM of the Gaussian filter in reciprocal space is shown as a red circle in the FFT (Fig. 4(d,f)). The radius of the Gaussian mask in nanometers is shown in real-time in the text box.
@@
where $N_{pix}$ is the number of pixels in one line, $L$ is the length of the image in nanometers, $v_{t}$ is the velocity (in nm/s) of the tip in one direction and the factor of 2 takes into account both the left and right fast-scan directions. This section is useful in estimating the time to record a “closed-feedback” $dI_{t}/dV_{b}$ map where the LDOS($eV_{b}$) is recorded at a single energy as the tip slowly scans the surface.
@@
The user inputs the time per spectra, which is the total time in seconds (including system overhead) to record a single spectrum, $T_{spec}$, i.e., one $dI_{t}/dV_{b}$ sweep. This section is useful in estimating the time to record an “open-feedback” LDOS($eV_{b}$) map where a $dI_{t}/dV_{b}(V_{b})$ spectrum is recorded at every pixel position. Together with the imaging window (Fig. 3(g)) provided by PyAtoms, this estimator is invaluable in ensuring that time-consuming QPI measurements will have the requisite reciprocal space resolution and spatial extent while avoiding aliasing effects.
@@ IV Applications
Owing to its simple and real-time GUI interface, PyAtoms is a useful research tool both during STM data collection as well as for data analysis. As shown in Fig. 2, PyAtoms can be used side-by-side during STM measurements in order to rapidly simulate experimental images and determine properties such as the constituent lattice constants and moiré twist angles between layers. Below we describe some examples of PyAtoms research capabilities.
@@
In practice, the strain tuning capabilities of PyAtoms can be used to simulate and determine either the amount of true, physical strain in STM images or the apparent strain due to scan artifacts such as tip drift. Figure 4 demonstrates the dramatic effect that strain can have on the moiré lattice and how PyAtoms can be used to determine this strain. Fig. 4(a) is a simulation of twisted trilayer graphene ($\theta_{12}=1.42^{\circ},\theta_{23}=-1.88^{\circ}$) moiré quasicrystal uri_superconductivity_2023 that features two moiré lattices. The atomic Bragg peaks for each graphene layer can be seen in the Fourier transform (Fig. 4(b)). The complex spatial structure of the moiré -of- moiré pattern (Fig. 4(c)) can be better seen after low-pass filtering to keep only the long wavelength periodicities (Fig. 4(d)). The presence of even miniscule strain in the top-most graphene layer ($\epsilon_{xx}=0.25$%) significantly alters the larger moiré pattern in both real (Fig. 4(e)) and reciprocal space (4(f)). For scanning probe researchers studying moiré multi-layers, PyAtoms can be used side-by-side during measurements to quickly determine the local strain in an area so that users can make informed decisions whether to seek other, less-strained areas.
Next, we show how PyAtoms can be used to determine and optimize scan parameters for performing QPI measurements, which reveal subtle scattering patterns at wavevectors that connect different portions of constant energy contours of the electronic band structure hoffman_imaging_2002; avraham_quasiparticle_2018. Here, we focus on NbSe2 which undergoes a strain-induced transition from an incommensurate $(3+\delta)\times(3+\delta)R0^{\circ}$ CDW to local areas of commensurate $(4\times 1)R0^{\circ}$ CDW order gao2018atomic. PyAtoms’ GUI allows users to view and optimize desired scan parameters to achieve high $k$-space resolution ($\delta k=2\pi/L$, image length $L$ large), sufficient $k$-space extent ($k_{max}=\frac{1}{2}N\delta k=\pi N/L=1.25|Q_{\text{Bragg}}|$, and image acquisition time, $T\approx 90$ min (Eq. 23), as shown in Fig. 5(a). Using these optimized parameters, the experimental real space image can be recorded at low tunneling bias (Fig. 5(b)) which, through its FFT (Fig. 5(c)) provides a QPI image near the Fermi energy joucken2020determination.
PyAtoms is additionally useful in simulating real-space CDW systems of various geometries and phases. Shown in Fig. 6 are experimental low-temperature STM topographic images of 2HH-NbSe2 with uniaxial strain applied along the vertical direction which has created a local transition from an incommensurate $(3+\delta)\times(3+\delta)R0^{\circ}$ CDW moncton_neutron_1977; soumyanarayanan2013quantum; arguello_visualizing_2014; gye2019topological to commensurate $(2\times 2)R0^{\circ}$ CDW order, in agreement with previous strain-STM experimentsgao2018atomic. The phase of the CDW in the STM images in Fig. 6(a)-(c) is strongly energy-dependent and strikingly similar to STM topographs of $1T$-TiSe2 pasztor_holographic_2019.
Figure 7: Site- and bond-centered charge density wave simulations. (a) Simulation of the approximately $(3\times 3)R0^{\circ}$ CDW in 2HH-NbSe2 ($a=3.44$ Å) which can be Se site- (left) or bond-centered (right). (b) “Star-of-David” $(\sqrt{13}\times\sqrt{13})R13.9^{\circ}$ CDW in 1TT-TaS2 ($a=3.36$ Å). (c) Kekulé-distorted graphene with $(\sqrt{3}\times\sqrt{3})R30^{\circ}$ bond-density wave texture which can be hollow- (left) or C site-centered (right). (d) Incommensurate CDW in quasi-1D K0.3MoO3, potassium blue bronze. (e)-(h) The respective Fourier transforms. The atomic peaks (yellow) and CDW peaks (blue) are highlighted. (a)-(b) simulated with “Simple” model ($\eta=0.3$); (c) and (d) simulated with “Log” model using $\xi=0$ and $\xi=1$, respectively.
PyAtoms can model a host of CDW orders in transition metal dichalcogenide (TMDC) van der Waals systems wilson_charge-density_1975. Figure 7(a,e) displays a simulation of 2HH-NbSe2, a TMDC that hosts a nearly commensurate CDW moncton_neutron_1977 with order that is approximately $(3\times 3)R0^{\circ}$ and is shown with a site-centered (left) and bond-centered (right) CDW phase. Figure 7(b,f) displays the $(\sqrt{13}\times\sqrt{13})R13.9^{\circ}$ CDW of TMDC $1T$-TaS2. Figure 7(c,g) displays simulations of graphene with a $(\sqrt{3}\times\sqrt{3})R30^{\circ}$ BDW, termed a Kekulé distortion with a phase centered at a hollow site (“Kekule-O,” left) li_scanning_2019; liu_visualizing_2022 or carbon site (“Kekulé-Y,” right) gutierrez_imaging_2016; nuckolls_quantum_2023. Finally, we show that PyAtoms can also be used to model the CDW in quasi-1D Peierls system K0.3MoO3, potassium blue bronze travaglini_blue_1981; brun_charge-density_2005. Fig. 7(d) shows the monoclinic atomic lattice projected onto the $(\bar{2}01)$ surface, which is what is measured by STM brun_charge-density_2005. The surface structure is simulated by a strained triangular lattice ($a_{1}=7.67$ Å, $\epsilon_{yy}=32\%$, image scan angle $\theta=90^{\circ}$). The nearly commensurate CDW, has a projection along the vertical chain direction (white arrow) ($|\mathbf{q}^{b}_{CDW}|/|\mathbf{b}^{*}|\approx 0.25$) for the ($\bar{2}01$) surface travaglini_blue_1981; brun_charge-density_2005 and is simulated with a strained square lattice ($b_{1}=1.39$ nm, $\epsilon_{xx}=\epsilon_{xy}=\epsilon_{yy}=18$%, $\theta_{12}=45^{\circ}$). This robust capability to simulate several families of CDW/BDW systems is especially useful when combined with the spectroscopic map time estimators (Fig. 3(i,j)) which allows users to efficiently plan detailed spectroscopic QPI measurements of quantum materials.
V Conclusion
@@
We have presented PyAtoms, an interactive open-source Python-based software for the real-time simulation of STM images of 2D materials. With its quick, intuitive, and robust simulation capabilities, we envision PyAtoms as a valuable companion during both STM measurements and STM data analysis. Future PyAtoms developments include offering additional methods for approximating images of multilayer systemsle2024moire, allowing generic one- and two-dimensional lattices, and, for the CDW/BDW simulations, full control of the phase in real space which would allow for improved simulation of measured STM data arguello_visualizing_2014; pasztor_holographic_2019; Shunsuke_PhysRevLett.132.056401. A current weakness in PyAtoms is that the comparison to experiment is largely qualitative. However, apart from its GUI, the open-source code that underlies PyAtoms can be called separately (e.g. through a Python terminal) and can be used, for example, to simulate moiré and superlattice systems beyond three layers. In principle, this ability to call on PyAtoms’ base functions offers the possibility to determine the best simulation parameters (lattice constant, A- and B-sublattice amplitude, strain) through a least-squares fitting to experimental STM data. This latter capability would elevate PyAtoms to a quantitative tool in the study of 2D materials. As an open-source software that allows for easy user-driven updates, we foresee PyAtoms becoming a valuable tool to the quantum materials scanning probe community.
Acknowledgements.
@@
The authors thank F. Joucken, K. Palotás for useful discussions and J. Berger, C. Irving for technical assistance with STM assembly. A.G.P acknowledges financial support from the National Science Foundation Graduate Research Fellowship Program (NSF-GRFP), the UCLA office of the Dean of Physical Sciences through the UCLA Physics Bridge Program, and the UCLA Graduate Division through a Eugene Cota-Robles Fellowship. K.-Y.W. acknowledges financial support from the Julian Schwinger Foundation. C.G. acknowledges teaching relief support from the UCLA offices of the Executive Vice Chancellor and Provost, Vice Chancellor for Research and Creative Activities, and the Vice Chancellor for Academic Personnel and partial support from the UCLA Hellman Society of Fellows. STM measurements were supported by UCLA Division of Physical Sciences startup funds. Uniaxial strain measurements were supported by a UBC-UCLA Collaborative Research Mobility Award via the Office of the Vice Chancellor for Research and Creative Activities at UCLA.
Author declarations
Conflict of Interest
The authors have no conflicts to disclose.
@@
C.G. created the original PyAtoms base codes (hexatoms, squareatoms, moirelattice) in Matlab. A.G.P. translated the base codes from Matlab to Python and created the GUI interface with input from C.G. Both A.G.P. and C.G. contributed to PyAtoms’ later updates. A.G.P., M.I.M.S., K.-Y.W., S.N., A.N., A.E.-I., and C.G. performed STM measurements. M.I.M.S., A.E.-I., and S.B. designed, assembled, and tested the STM uniaxial strain device. C.G. wrote the manuscript and created the figures with input from A.G.P and M.I.M.S. All authors contributed to editing the manuscript.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.